OpenStudy (gatsbyyy):

3/10x-6/11=7/10x-9/11

OpenStudy (prizzyjade):

\[\frac{ 3 }{ 10x} - \frac{ 6 }{ 11 } = \frac{ 7 }{ 10x } -\frac{ 9 }{ 11}\]

OpenStudy (prizzyjade):

is that the right equation?

OpenStudy (gatsbyyy):

No, i'll make a drawing of it.

OpenStudy (prizzyjade):

Hahaha ok

OpenStudy (prizzyjade):

\[\frac{ 3 }{ 10 }x-\frac{ 6 }{ 11 }=\frac{ 7 }{ 10 }x-\frac{ 9 }{ 11 }\]

OpenStudy (prizzyjade):

use the transposition method to combine the fractions with the same decimal

OpenStudy (gatsbyyy):

What's the transposition method?

OpenStudy (thecatman):

are you solving for x

OpenStudy (gatsbyyy):

Yes

OpenStudy (thecatman):

one moment

OpenStudy (gatsbyyy):

kk

OpenStudy (prizzyjade):

since your finding x combine the two expression with x in it

OpenStudy (thecatman):

combine factions

OpenStudy (gatsbyyy):

How do I do that?

OpenStudy (thecatman):

3/10 x and 7/10 x =1x

OpenStudy (gatsbyyy):

Ohh Okay

OpenStudy (gatsbyyy):

So far I have: \[-\frac{ 6 }{ 11 }=x-\frac{ 9 }{ 11}\]

OpenStudy (thecatman):

sorry fing the gcf

OpenStudy (gatsbyyy):

I got 3/11 as my answer

OpenStudy (thecatman):

33/110 x - 60/110 = 77/110x - 90/110

OpenStudy (thecatman):

now try

OpenStudy (gatsbyyy):

Why would i have to do that if I got rid of the 10 fractions since the 3/10 + 7/10 adds up to just 1?

OpenStudy (thecatman):

math is complicated but this is how my teacher taught it